SPS LS1 Planning meeting (No.26) 11/12/13 - PowerPoint PPT Presentation

paul2
sps ls1 planning meeting no 26 11 12 13 n.
Skip this Video
Loading SlideShow in 5 Seconds..
SPS LS1 Planning meeting (No.26) 11/12/13 PowerPoint Presentation
Download Presentation
SPS LS1 Planning meeting (No.26) 11/12/13

play fullscreen
1 / 20
Download Presentation
SPS LS1 Planning meeting (No.26) 11/12/13
270 Views
Download Presentation

SPS LS1 Planning meeting (No.26) 11/12/13

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. SPS LS1Planning meeting (No.26)11/12/13 David Mcfarlane SPS.Technical-Coordination@Cern.ch

  2. Topics • Lifts and Monte-charges. • Upcoming works. • Maintenance 18 kV. • CSA presentation • Planning. • False floors open in the Bas. • E.C.Rs. • Christmas Vacation • Access rights and safety equipment. • Important Information. • A.O.B SPS.Technical-Coordination@Cern.ch

  3. Lift Maintenance Dec/Jan SPS.Technical-Coordination@Cern.ch

  4. Upcoming works • Replacement of Irradiated cables (BA1) • Now until March 2014 • Civil Engineering works (TT10) • Now until End of year • Replacement of Septas (BA2) • 28th Oct until end of year • Kicker conditioning (BA6) • Dec to mid Jan • Re-installation of beam line and equipment in TT10 • January 2014 SPS.Technical-Coordination@Cern.ch

  5. Maintenance 18 kV EMD2*9 JURA du vendredi 20 dec au samedi 21 dec. 2013 La maintenance 18 kV  de la station majeure de JURA ME9 est impérative et critique La maintenance des barres des tranches 1 et 3  en cours actuellement . Elle n’affecte pas le fonctionnement de l’auto transfert et du réseau secours.  Au contraire de la maintenance des barres de la tranche 2 qui nécessitera la mise hors service du système auto transfert et de la liaison Normal/secours. L’absence de la fonction  secours ME9 a pour conséquence l’interdiction d’accès en souterrain LHC1 –ATLAS –LHC18 – SPS6 –SPS7 pendant les 3 à 4  jours de maintenances Par ailleurs des interventions du contractant Schneider E078 sont déjà planifiées W50 et W51 au LHC6 et LHC7… Suite à nos divers échanges de ce jour et au TIOC, nous pouvons proposer le planning suivant Lundi 16 déc. au mercredi 18 déc. : maintenance 18kv LHC7 selon planning LHC Jeudi 19 déc. : jour de repos des techniciens Schneider (et OP) Vendredi 20 déc. au dimanche 22 déc. : maintenance des barres de la tranche 2 JURA : -> pas d’accès en souterrain LHC1 –ATLAS –LHC18 – SPS6 –SPS7 A noter qu’il n’y aura pas de coupure utilisateur, les réseaux aval étant réalimentés sans coupure ACCESS INTO BA6 &BA7 WILL BE BLOCKED FROM THURSDAY LUNCHTIME (19TH DECEMBER) SPS.Technical-Coordination@Cern.ch

  6. CSA presentation SPS.Technical-Coordination@Cern.ch

  7. NOTRE MISSION • Interface avec les responsables des accès des zones (GS-ASE) • S’assurer que les accès aux accélérateurs sont fermés, • Faciliter l’accès des usagers/intervenants, • S’assurer que les intervenants respectent les réglements • Droits d’accès, • Port des EPI, • Bonne utilisation PAD/MAD • intervenir sur les alarmes • Faire remonter les anomalies liées aux accès • Vers TI (problème technique) • Vers GS-DI (non-conformité des personnels / non respect des procédures)

  8. NOTRE BESOIN • Échanciersdes interventions par zone • Appeler l’agent «ZORA» (n° 6 25 11) • En début et fin d’intervention • Quand une porte est «forcée» dans les accélérateurs • Refermer toutes les portes que vous ouvrez • Autres problèmes lié aux accès

  9. SPS LS1 schedule Open Access : 27/03/13 Close Access : 27/06/14 Beam off : 16/02/13 Beam on: 13/10/14 Beam on Cool down/RP survey and Magnet tests Cool down NO ACCESS Kicker Survey General Maintenance Survey General Maintenance Survey Survey Cables EPC works RF upgrade works RF upgrade works EPC works Survey Cables Survey Kicker Survey 18 KV Civil Engineering in TT10 Fibres Magnets Magnets Magnets Magnets Magnets Magnets Kicker Cables Magnets Magnets 18 KV General maintenance General Maintenance Survey General maintenance CrabCavities General Maintenance Cooling tower 863 Vacuum Septa works Septa works Vacuum Vacuum Ring-main Ring-main Ring-main Ring-main Ring-main Fibres Ring-main Ring-main General Maintenance 18 KV General Maintenance Vacuum Prep work in LSS1 Vacuum Survey Horizontal Vertical smoothing RF Upgrade works Cables Cool down Limited ACCESS UA9 works TT20 pipework 18 KV Fibres CV maintenance CV maintenance Civil Engineering CV maintenance 18 KV Fibres Irradiated Cables Cables Cables Cables Fibres Cables 18 KV EL Studies Civil Engineering TT10 Consolidation CV studies Cables Septa replacement Civil Engineering Civil Engineering Civil Engineering CrabCavities RF Upgrade works Civil Engineering RF Upgrade works Re-install equipment removed for cable works RAL CV works Magnets Fibres Cables Magnets Magnets Magnets Magnets Magnets Magnets Magnets EPC Commissioning tests Vacuum DSO tests/HV tests Cold check out & Start-up Protons to the North Area Protons to LHC Ions to NA61 CV Controls upgrade Re-route Cabling

  10. False floors lifted in Surface BAs • Below is a table showing the dates that EL will be pulling cables at each point of the SPS. • During these periods large sections of the false floor in these surface buildings will be lifted. • Any works in the BA’s that involve lifting of the false floors MUST be reported to the Coordination team • If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first. SPS.Technical-Coordination@Cern.ch

  11. Engineering Change Requests • Templates and Guidelines for Engineering Change Requests: • EDMS: 1271880 • SPS ECR template (in DFS): • https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/Applications/CERN/Workgroup%20Templates/SPS/ECR-Template-SPS.dotx SPS.Technical-Coordination@Cern.ch

  12. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Removal of AEWB Pick-up from position 31932 to 31502 • (EDMS No. 1245361) RELEASED • Removal of a BPWA Pick-up from position 31931 to 31101 • (EDMS No. 1245393)RELEASED • Installation of a new Wideband Kickers system in position 32173 • (EDMS No. 1245400)RELEASED • Installation of a new HBW Damper Pick-up in place of BPWA 31931 & AEWB 31932 • (EDMS No. 1245401)RELEASED • Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 • (EDMS No. 1246559)RELEASED • Reinstallation of pump and gauges (Pirani + Penning) in 11993 (vacuum sector 150) • (EDMS No. 1249305)RELEASED • Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) • (EDMS No. 1249307 v.2) RELEASED • Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. • (EDMS No. 1075945) UNDER IMPLEMENTATION SPS.Technical-Coordination@Cern.ch

  13. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. • (EDMS No. 1221341) RELEASED • SPS Beam Position Monitoring Prototyping Setup in LSS4. • (EDMS No. 1099630)RELEASED • Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. • (EDMS No. 1251693)RELEASED • Installation d'un nouveau moniteur de position de type BPCR en LSS2 au 22105. • (EDMS No. 1251695) RELEASED • Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. • (EDMS No. 1253832) UNDER IMPLEMENTATION • Interlocking of Injection Tunnel upstream areas with LHC powering. • (EDMS No. 1257949)RELEASED • Move of BPH.31998 to BPH.42173. • (EDMS No. 1272718) RELEASED • Optical Fibre Racks. • (EDMS No. 1273766)UNDER IMPLEMENTATION SPS.Technical-Coordination@Cern.ch

  14. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Removal of old fixed TMP vacuum groups in 11655-11691. • (EDMS No.1273232)RELEASED • BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). • (EDMS No. 1282651) UNDER IMPLEMENTATION • Installation de Coffrets d'Alimentation en LSS1 et TT10. • (EDMS No. 1282860) RELEASED • Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. • (EDMS No. 1288607)UNDER IMPLEMENTATION • Installation de nouveaux chemins de câbles en LSS1+ du SPS. • (EDMS No. 1291889 v.2) RELEASED • TPSN replacement in LSS2 of SPS. • (EDMS No. 1302672 v.1) RELEASED • Installation d'un nouveau chemin de câbles en LSS2- du SPS. • (EDMS No. 1325439 v. 0.1) IN WORK • Replacement of pumps in 3 tanks MKP in LSS1 of the SPS . • (EDMS No. 1328248 v. 0.1)UNDER APPROVAL SPS.Technical-Coordination@Cern.ch

  15. Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN-0000076703:V0/P:CERN-0000096067:V0/TAB3 • Suppression of VVFA 11899 in LSS1 of SPS. • (EDMS No. 1328249v. 0.1) UNDER APPROVAL • New shielding wall behind TIDV 11892 in SPS . • (EDMS No. 1331621v. 0.1)IN WORK SPS.Technical-Coordination@Cern.ch

  16. The Christmas Vacation • There is no planned works or accesses into the SPS during the 2 week vacation. • There are no requests for services during this period?? • BA6 and BA7 will be closed from 12:00 on Thursday the 19th December SPS.Technical-Coordination@Cern.ch

  17. Access rights and safety equipment • Everyone who enters the SPS MUST have the following: • Completed the correct courses. • Safety course levels 1,2 and 3. • HabilitationElectrique. • Radiological Protection. • Self Rescue Mask Training. • The correct access authorisation (via EDH). • A valid approved IMPACT. (with their name included on this IMPACT) • They must know the number of this IMPACT. • Safety shoes. • Biocell. • Electronic dosimetre. • Passive dosimetre. • Helmet plus Light. • Please note that personnel are not allowed to work in the tunnel alone!! SPS.Technical-Coordination@Cern.ch

  18. Important information • Useful Contact Numbers • SPS Coordination Team • David MCFARLANE- 164247 • Franck BAIS – 165467 • James RIDEWOOD - 160590 • Control Room • SPS CCC – 77500 • TI CCC – 72201 • CSA - 62511 • RP • Christophe Tromel – 163199 • Angelito Herve – 163168 • Nadine Conan (BA1) – 160641 • Elodie Aubert - 169097 • Kurt Weiss (BA1) -160759 • Coordination Safety • Emmanuel Paulat – 163870 • John Etheridge - 164647 SPS.Technical-Coordination@Cern.ch

  19. A.O.B. • Please make sure you fill out you IMPACTS correctly!!!! • There are currently 19 IMPACTs that are “LATE”. If you have completed your work then you need to close the related IMPACT. • If you have ANY concerns about not being complete before the 27th June 2014. I need to know about it NOW!!!! • The SPS cooling towers will be stopped 03/02/14 to 07/02/14 for EL works on switchboard, transformer and SPS pumping station • LS2!!!!!! Work has already begun on the planning for LS2. If you are planning on doing any major works in LS2 for consolidation then we need to know about it now. • Please can you send me an e-mail stating the main tasks you are planning/hoping to do during LS2 SPS.Technical-Coordination@Cern.ch

  20. SPS coordination website http://sps.web.cern.ch/SPS/ http://cern.ch/LS1planning \\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 08/01/2014 112-R-028 SPS.Technical-Coordination@Cern.ch